OpenStudy (anonymous):

Solve 2x + y = 7 x – y = –4

OpenStudy (anonymous):

Add the first and second equations. The result will be a linear equation in x; solve that equation. Take the solution for x, substitute to get the value for y.

OpenStudy (anonymous):

x=1,y=5

OpenStudy (anonymous):

Thanks guys.

OpenStudy (anonymous):

\[(2x + y = 7)+( x – y = –4) \implies x=1 \implies (x,y)=(1,5)\]The first bit of notation isn't a thing of beauty, but I think you can get the idea.